Sheridan Company had 808000 shares of common stock outstanding at December 31, 2018. In addition, it had 159000 stock options outstanding, which had been granted to certain executives, and which gave them the right to purchase shares of Sheridan's stock at an option price of $37 per share. The average market price of Sheridan's common stock for 2018 was $50. What is the number of shares that should be used in computing diluted earnings per share for the year ended December 31, 2018?

925660
849340
808000
863865
Answer #1

Calculate number of shares used in diluted earning per share

Stock option diluted = 159000-(159000*37/50) = 41340

Number of share used in diluted earning per share = 808000+41340 = 849340

SO answer is b) 849340